OpenStudy (anonymous):

if |a|,|b|<1, then find the sum of the series:- a(a+b) + a^2(a^2+b^2) +a^3(a^3+b^3) + ........infinity. Note : this is not in GP series. Then how can we find the sum of series

OpenStudy (anonymous):

@ExperimentX

OpenStudy (anonymous):

plse help me

OpenStudy (anonymous):

@Shubhamsrg Plse help me

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

plse help me

OpenStudy (shubhamsrg):

well this is not in GP series but you can convert it into 2 separate GPs why dont you expand all the brackets ?

OpenStudy (anonymous):

how

OpenStudy (shubhamsrg):

a(a+b) = a^2 + ab a^2 (a^2 + b^2 ) = a^4 + a^2 b^2 like that.

OpenStudy (shubhamsrg):

following?

OpenStudy (anonymous):

wait ..i m trying

OpenStudy (anonymous):

yes ..i m getting sam e vat u have written , vat after that

OpenStudy (shubhamsrg):

Why didnt you simply write it here? What was the need for a doc file ? o.O

OpenStudy (shubhamsrg):

anyways, try forming 2 separate GPs 1 would be a^2+ a^4 + a^6 .... can you tell me second ?

OpenStudy (anonymous):

ab+a^2b^2+a^3b^3

OpenStudy (anonymous):

but we will take common ab from second

OpenStudy (anonymous):

ab(1+ab+a^2b^2)

OpenStudy (shubhamsrg):

you may, though there is no need..

OpenStudy (anonymous):

k

OpenStudy (shubhamsrg):

so now you have 2 GPs, both infinite and common ratio < 1 ..so you can do further right ?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

on e thing more..it means we will be getting two G.P. + two answers.

OpenStudy (anonymous):

i m getting answer, while solving first G.P a^2/(1-a^2) and after solving second one , i get ab/(1-ab)

OpenStudy (shubhamsrg):

so final answer will be sum of both those sums, ofcorse.

OpenStudy (anonymous):

[a^2/(1-a^2)]+[ab/(1-ab)]

OpenStudy (shubhamsrg):

yep.

OpenStudy (anonymous):

thanks again